# Gamma/hadron discrimination through the analysis of the shower footprint at low energies

### v1

@@ Abstract
Gamma/hadron discrimination in ground-based gamma-ray observatories at the GeV energy range is challenging as traditional muon-based methods become less effective at lower energies. This work explores a novel gamma/hadron discrimination method for Extensive Air Shower arrays that analyzes the shower signal footprint patterns using advanced Vision Transformer (ViT) models. The robustness of this method against noise, including atmospheric muons and low-energy proton showers, demonstrates its potential for application in future gamma-ray observatories.
@@ I Introduction
The study of gamma rays is crucial for understanding extreme astrophysical events and probing fundamental physics. Gamma-rays in the sub-TeV to TeV range can provide insights into Active Galactic Nuclei (AGNs), Gamma Ray Bursts (GRBs), and potential new physics beyond the standard model, including dark matter research De Angelis and Pimenta (2018); Viana et al. (2019).
Their detection above a few hundred of GeV is possible using ground-based detectors as satellite detection is limited by low fluxes Atwood et al. (2009). These experiments indirectly detect gamma rays at very high altitudes by using Extensive Air Showers (EAS), which are a cascade of secondary shower particles produced in the atmosphere due to the interaction of the gamma-ray with it. Then, the characteristics of the primary particle such as energy and direction are inferred from the secondary particles detected at the ground. This method enables surveying large portions of the sky and is sensitive to transient phenomena. However, gamma rays must be distinguished from the dominant hadronic background Degrange and Fontaine (2015); Albert et al. (2019); Torres Peralta et al. (2024).
Muon identification remains one of the main methods for gamma/hadron separation at TeV energies Conceição et al. (2021); González et al. (2022, 2020); Kunwar et al. (2023). However, the muon content of hadronic showers is scarce at sub-TeV energies. Alternatively, one may analyze the shower footprint patterns on the ground to infer features of the shower development Iacovacci et al. (2013); Conceição et al. (2022, 2024); Abeysekara et al. (2017). Hadron-induced showers, unlike pure electromagnetic showers, generate particles with high transverse momentum, causing the shower to spread more laterally and form clusters. Initial studies have shown great potential in using these ground patterns for gamma/hadron discrimination, but further work is needed to fully understand and optimize their effectiveness against noise Assunção et al. (2019).
This study demonstrates the potential of employing advanced Vision Transformer (ViT) models Vaswani et al. (2017); Dosovitskiy et al. (2020) to accurately discriminate between gamma and hadron-induced air showers. The footprint image, created from signals detected by the individual Water Cherenkov Detectors (WCDs) in the detector array, is used as input for the model. The method’s robustness against noise, including atmospheric muons and low-energy proton showers, is also assessed.
@@ II Set of simulations
The Extensive Air Showers were simulated with CORSIKA (version 7.5600) Heck et al. (1998), taking FLUKA Ferrari et al. (2005); Böhlen et al. (2014) and QGSJet-II.04 Ostapchenko (2011) as the low- and high-energy hadronic interaction model, respectively. The detector response was emulated by summing the expected signals from particles hitting the station. These signals were derived using specific parameterizations based on the energy of protons, muons, and electrons/gammas, created by injecting uniformly sampled vertical particles onto a small water Cherenkov detector (WCD) station with three PMTs at the bottom Assis et al. (2022). These parameterizations included both the mean signal and its standard deviation, allowing us to simulate fluctuations in the WCD signal response due to particle interactions and light collection. Additional fluctuations for muons, taking into account the variations in the tracklength caused by trajectory changes, were included using muon distribution data from proton-induced shower simulations with Geant4 (version 4.10.05.p01) Agostinelli et al. (2003); Allison et al. (2006, 2016).
@@
The CORSIKA simulations were generated using an $E^{-1}$ spectrum to balance computational time with statistical accuracy. During the analysis, weights of $E_{0}^{-1}$ and $E_{0}^{-2}$ were applied to gammas and protons respectively to ensure a realistic power-law energy spectrum. The showers are nearly vertical, with simulated zenith angles of $\theta_{0}=10^{\circ}$ and $\theta_{0}\in\left[5^{\circ};10^{\circ}\right]$ for gammas and protons, respectively.
Each CORSIKA shower was simulated six times with different core positions to enhance statistical reliability. The simulated cores where distributed uniformly over the array, and its impact on the gamma/hadron discrimination power is assessed in section IV.2. The event was simulated in four conditions: without noise, with noise from atmospheric muons, noise from low-energy proton showers, and noise from both sources. The dataset is split such that a given CORSIKA shower appears in only one dataset, regardless of its core position or contamination. This approach ensures realistic results, preventing data leakage, where the model could learn a specific CORSIKA shower in training and test it correctly in other sets. It also helps the network learn that the CORSIKA shower corresponds to a given primary particle, independent of contamination or core position.
[…]
The number of events generated after having simulated six core positions for each CORSIKA shower and added the above-mentioned contamination sources are detailed for each energy bin in Table 1.
Energy Split Gammas Protons Total
200 GeV Train 65 6176561765\,61765 617 20 8312083120\,83120 831 86 4488644886\,44886 448
Test 17 5421754217\,54217 542 6 41964196\,4196 419 23 9612396123\,96123 961
Validation 10 2371023710\,23710 237 3 24232423\,2423 242 13 4791347913\,47913 479
500 GeV Train 58 1605816058\,16058 160 45 3084530845\,30845 308 103 468103468103\,468103 468
Test 16 3691636916\,36916 369 13 9161391613\,91613 916 30 2853028530\,28530 285
Validation 9 57295729\,5729 572 9 21692169\,2169 216 18 7881878818\,78818 788
1 TeV Train 39 0083900839\,00839 008 31 8333183331\,83331 833 70 8417084170\,84170 841
Test 9 40394039\,4039 403 11 7681176811\,76811 768 21 1712117121\,17121 171
Validation 9 03490349\,0349 034 5 25852585\,2585 258 14 2921429214\,29214 292
@@ 1 TeV Train 39 0083900839\,00839 008 31 8333183331\,83331 833 70 8417084170\,84170 841
Figure 1: Energy calibration using gamma-induced showers. The black line represents the calibration and the red points the mean signal for each energy bin.
Figure 2: Bias and resolution of the energy calibration. The grey area highlights the region with a bias of $\leq 5\%$. The resolution is fitted with the formula: $y=\sqrt{a/\sqrt{x}+b}$.
@@ III Analysis strategy: Vision Transformers
ViTs Vaswani et al. (2017); Dosovitskiy et al. (2020) are emerging in the field of computer vision. Unlike traditional Convolutional Neural Networks (CNNs) O’Shea and Nash (2015), which rely on convolutional and fully connected layers, ViTs depend on self-attention mechanisms, enabling them to capture long-range dependencies and context within images. To do this, the image is divided into fixed-size patches, each of which is linearly embedded, followed by the addition of position embeddings. Subsequently, the resulting sequence of vectors is fed into a standard Transformer encoder, where extra learnable classification tokens are added for classification Dosovitskiy et al. (2020). This innovation enables these models to capture intricate long-range dependencies and contextual information within images, leading to enhanced capabilities in various applications.
@@
The pre-trained ViT model “google/vit-base-patch16-224-in21kDosovitskiy et al. (2020, 2021) has been employed in this study. This is a transformer encoder model (BERT-like) pre-trained on a large data set called ImageNet-21k Ridnik et al. (2021), which is composed of 14 million images at a resolution of 224x224 pixels and has 21 8432184321\,84321 843 classes. Afterwards, a linear layer was placed on top of the final hidden state of the [CLS] token and trained with our problem’s data for 30 additional epochs with a batch size of 32, the default learning rate of $5\cdot 10^{-5}$ and warm-up ratio of $10\%$. This pipeline was done with the Hugging Face Transformers open-source library Wolf et al. (2019), which provides APIs and tools to download and train state-of-the-art pre-trained models.
The input of the model is an image created by mapping the positions of the WCDs and assigning colours based on the total signal in each tank. To emphasize signal structures within the footprint and group stations with similar signals, a linear contour plot with 5 levels and a grid with cells of similar size to the stations is overlaid. A logarithmic scale is applied to prevent saturation of the colour scale due to the high signals in stations near the shower core. To reduce noise from background or late-arriving particles, the shower plane is fitted using the positions of all active stations, with the time of the first particle arrival at each station as the Z coordinate Apolinário et al. (2023). It was found that with an average background of 23 stations, the plane fit successfully converges when at least 30 stations are present in the event. Therefore, only events with more than 30 active stations (signal + background) were considered for analysis. Following this procedure, only stations with a temporal offset of less than 505050\,50ns from the fitted shower plane were selected.
@@
Finally, the objective function of the method is defined as the classification probability, denoted as $P\in\left[0;1\right]$, that a proton has induced the analyzed shower.
IV Results
In this section, the gamma/hadron separation capability of the method is evaluated. The background rejection factor is computed for a selection efficiency of roughly $75\%$ and compared with the results achieved by the LATTES Assis et al. (2018), HAWC Albert et al. (2024), and LHAASO-WCDA Wang et al. (2019); Aharonian et al. (2021) experiments. The background rejection factor for a gamma efficiency selected to achieve the optimal Q-factor is also presented in Appendix V.1. A single model is trained for each energy bin and assessed under various noise conditions.
@@ IV.1 Resilience to Noise from Atmospheric Muons and Low-Energy Proton Showers
Figure 5: Proton rejection efficiency at different energies and noise sources.
Figure 6: Dependence of the efficiency on the shower core position at different energy bins for noise-free showers.
For an EAS array with the given characteristics, the number of stations triggered by atmospheric muons within a 200200200\,200ns time window is expected to follow a Poisson distribution with a mean of 23 stations Apolinário et al. (2023). To assess the resilience of the method against noise from atmospheric muons, each shower event is contaminated with a random number of stations based on this distribution. The muonic signal added to the contaminated tanks is randomly selected using a probability function derived from the signal distribution of approximately 150 000150000150\,000150 000 single muon events from proton showers in the $\sim 1\,\rm TeV$ energy bin.
@@
For each of the selected reconstructed energy bins, a ViT model is trained and tested using the original shower-derived images and their two extra images with contamination. The results for the test data set are shown in Figure 5. The results show a minimal impact of adding noise from low-energy proton showers on the gamma/hadron separation method, with negligible differences observed across energy bins. However, the noise from atmospheric muons does have a more noticeable effect as the energy increases. This trend is expected to be related to the fact that low-energy showers rarely contain muons, leading the model to rely less on those features for classification. The same behaviour was found for the worst-case scenario in which both noise sources are superimposed in the shower footprints. The developed model was able to outperform LATTES, HAWC and LHAASO across all tested conditions.
IV.2 Dependence on the position of the shower core
To assess the model performance dependence on the position of the shower core, two subsets of noise-free shower events were created according on their true core position: $r_{\text{core}}\in\left[0,50\right]\,\text{m}$ and $r_{\text{core}}\in\left[100,150\right]\,\text{m}$. The same models used in the previous subsection were employed for this analysis.
The results, shown in Figure 6, indicate that although there is a slight degradation in performance, particularly at low energies where the shower footprint is smaller, the difference is minimal between events with simulated shower cores near the array center and those farther away. Thus, a limited dependence on the shower core position has been observed.
IV.3 Dependence on the muon content of the shower

### v2

# Discriminating sub-TeV gamma and hadron-induced showers through their footprints

@@ Abstract
Gamma/hadron discrimination in ground-based gamma-ray observatories at the sub-TeV energy range is challenging as traditional muon-based methods become less effective at lower energies. This work explores a novel gamma/hadron discrimination method for Extensive Air Shower arrays that analyzes the shower signal footprint patterns using a state-of-the-art pre-trained Vision Transformer (ViT). The resilience of the method to background noise, such as atmospheric muons and low-energy proton showers, along with its adaptability to different zenith angles and array configurations, demonstrates its potential for application in current and future ground-based gamma-ray observatories.
@@ I Introduction
The study of gamma rays is crucial for understanding extreme astrophysical events and probing fundamental physics. Gamma-rays in the sub-TeV to TeV range can provide insights into Active Galactic Nuclei (AGNs), Gamma Ray Bursts (GRBs), and potential new physics beyond the standard model, including dark matter research [1, 2].
Their detection above a few hundred of GeV is possible using ground-based detectors as satellite detection is limited by low fluxes [3]. These experiments indirectly detect gamma rays at very high altitudes by using Extensive Air Showers (EAS), which are a cascade of secondary shower particles produced in the atmosphere due to the interaction of the gamma-ray with it. Then, the characteristics of the primary particle such as energy and direction are inferred from the secondary particles detected at the ground. This method enables surveying large portions of the sky and is sensitive to transient phenomena. However, gamma rays must be distinguished from the dominant hadronic background [4, 5, 6].
Muon identification remains one of the main methods for gamma/hadron separation at TeV energies [7, 8, 9, 10, 11]. However, the muon content of hadronic showers is scarce at sub-TeV energies. Alternatively, one may analyze the shower footprint patterns on the ground to infer features of the shower development [12, 13, 14, 15]. Hadron-induced showers, unlike pure electromagnetic showers, generate particles with high transverse momentum, causing the shower to spread more laterally and form clusters. Initial studies have shown great potential in using these ground patterns for gamma/hadron discrimination, but further work is needed to fully understand and optimize their effectiveness against noise [16, 17].
Machine learning (ML) has recently become an innovative instrument in the field of physics, especially within astroparticle physics, enabling advancements in numerous domains that require extensive data analysis. In modern cosmic and gamma-ray observatories, ML techniques have shown significant potential to enhance gamma/hadron separation [18, 19, 20, 21, 22, 23, 24], event reconstruction [25, 26, 27, 28], and even allow neutrino identification with water Cherenkov detectors (WCDs) [29]. These breakthroughs are analogous to significant advances in related domains, such as the IceCube detection of neutrinos from the galactic plane [30], investigations on the mass composition and $X_{\rm max}$ estimation of ultra-high-energy cosmic rays (UHECRs) by the Pierre Auger Observatory [31, 32, 33], and the search of new physics with the Large Hadron Collider (LHC) [34, 35, 36, 37, 38].
Among ML-based techniques, transformers, which use attention mechanisms for data analysis [39], have emerged as particularly effective. Their initial applications in event reconstruction for astroparticle experiments have demonstrated promising results [40]. In this work, we demonstrate that state-of-the-art pre-trained Vision Transformers (ViTs) [41] have significant potential to accurately discriminate between gamma and hadron-induced air showers. The footprint image, created from signals detected by the individual WCDs in the detector array, is used as input for the model. The method’s robustness against noise, including atmospheric muons and low-energy proton showers, is also assessed.
@@ II Set of simulations
Following the same simulation strategy as in [8, 7], the Extensive Air Showers were simulated with CORSIKA (version 7.5600) [42], taking FLUKA [43, 44] and QGSJet-II.04 [45] as the low- and high-energy hadronic interaction model, respectively. The detector response was simulated with the Geant4 toolkit (version 4.10.05.p01) [46, 47, 48].
The experimental observation level was established at 5 20052005\,2005 200 m above sea level  [^1] This altitude corresponds to the altitude of the ALMA site in Chile.. The detector array comprises 5 72057205\,7205 720 Mercedes WCDs [7], covering an area of 80 $80\,000\,\rm m^{2}$ with a fill factor of approximately $85\%$. Two additional arrays, with fill factors of $65\%$ and $40\%$, were generated from the initial simulation through the uniform removal of stations, as explained in Subsection IV.4 (see Figure 8).
The CORSIKA simulations were generated using a $E^{-1}$ spectrum to balance computational time with statistical precision. During the analysis, event weights were applied to correct the spectra of gamma rays and protons to spectral indices of $E^{-2}$ and $E^{-3}$, respectively, to ensure a realistic power-law energy spectrum of cosmic rays.
The showers are nearly vertical, with simulated zenith angles of $\theta_{0}=10^{\circ}$ and $\theta_{0}\in\left[5^{\circ};15^{\circ}\right]$ for gammas and protons, respectively. To evaluate the effectiveness of the method for showers with greater inclination, simulations were also conducted for gamma and proton events at zenith angles of $\theta_{0}=30^{\circ}$ and within the range $\theta_{0}\in\left[25^{\circ};35^{\circ}\right]$, respectively. The choice of these wider zenith angle ranges for protons reflects the diffuse nature of the background, as a realistic scenario would involve direction reconstruction and accept events with slightly different zenith angles. Nonetheless, as discussed in Subsection IV.2, the performance of the ViT exhibits minimal dependence on the zenith angle.
Each CORSIKA shower was simulated five times using a Geant4 dedicated simulation with different core positions to enhance statistical reliability. The simulated cores were uniformly distributed within a radius of $165\,\rm m$, slightly exceeding the radius of the array of $160\,\rm m$, and their impact on the gamma/hadron discrimination power is assessed in Subsection IV.3. The event was simulated under four conditions: without noise, with noise from atmospheric muons, noise from low-energy proton showers, and noise from both sources. The data set is split such that a given shower appears in only one dataset, regardless of its core position or contamination. This approach ensures realistic results, preventing data leakage, where the model could learn a specific shower in training and test it correctly in other sets. It also helps the network learn that the shower corresponds to a given primary particle, regardless of contamination or core position.
@@
The number of events generated after having simulated five core positions for each shower and adding the above-mentioned contamination sources are detailed for each energy bin in Table 1. The data set was divided into training, validation, and test sets, where the training set is used to train the model, the validation set is used to tune hyperparameters and assess performance during development, and the test set is reserved for evaluating the final model performance.
Energy $\bm{\theta}$ Split Gammas Protons Total
200 GeV $10^{\circ}$ Train 79 7477974779\,74779 747 77 3577735777\,35777 357 157 104157104157\,104157 104
$10^{\circ}$ Test 22 5822258222\,58222 582 22 7912279122\,79122 791 45 3734537345\,37345 373
$10^{\circ}$ Validation 13 7871378713\,78713 787 12 7031270312\,70312 703 26 4902649026\,49026 490
$30^{\circ}$ Test 1 56315631\,5631 563 1 28612861\,2861 286 2 84928492\,8492 849
500 GeV $10^{\circ}$ Train 54 9475494754\,94754 947 57 0225702257\,02257 022 111 969111969111\,969111 969
$10^{\circ}$ Test 15 7901579015\,79015 790 16 5251652516\,52516 525 32 3153231532\,31532 315
$10^{\circ}$ Validation 8 88388838\,8838 883 9 50295029\,5029 502 18 3851838518\,38518 385
$30^{\circ}$ Test 1 20812081\,2081 208 $968$ 2 17621762\,1762 176
1 TeV $10^{\circ}$ Train 43 6694366943\,66943 669 44 5324453244\,53244 532 88 2018820188\,20188 201
$10^{\circ}$ Test 12 4471244712\,44712 447 12 6121261212\,61212 612 25 0592505925\,05925 059
$10^{\circ}$ Validation 7 03070307\,0307 030 7 39473947\,3947 394 14 4241442414\,42414 424
$30^{\circ}$ Test $899$ $584$ 1 48314831\,4831 483
@@ 1 TeV $10^{\circ}$ Train 43 6694366943\,66943 669 44 5324453244\,53244 532 88 2018820188\,20188 201
Figure 1: Energy calibration using gamma-induced showers with $\theta_{0}=10^{\circ}$ and $r_{core}\in\left[0;165\right]\,\rm m$. The black line represents the calibration and the red points the mean signal for each energy bin.
Figure 2: Bias and resolution of the energy calibration. The grey area highlights the region with a bias of $\leq 5\%$. A smooth curve was incorporated to guide the eye.
@@ III Analysis strategy: Vision Transformers
ViTs are emerging in the field of computer vision. Unlike traditional Convolutional Neural Networks (CNNs) [50], which rely on convolutional and fully connected layers, ViTs depend on self-attention mechanisms, enabling them to capture long-range dependencies and context within images [39]. To do this, the image is divided into fixed-size non-overlapping patches, each of which is linearly embedded, followed by the addition of position embeddings. Subsequently, the resulting sequence of vectors is fed into a standard Transformer encoder, where extra learnable classification tokens are added for classification. This innovation enables these models to capture intricate long-range dependencies and contextual information within images, leading to enhanced capabilities in various applications.
In this study, the state-of-the-art pre-trained ViT model for classification “google/vit-base-patch16-224-in21k[41, 51, 52, 53] has been used. This model is a BERT-like transformer encoder that includes an extra linear layer applied to the classification token ([CLS]) for executing classification tasks. It was pre-trained on a large data set called ImageNet-21k [54], which is composed of 14 million images at a resolution of $224\times 224$ pixels and has 21 8432184321\,84321 843 classes. It is crucial to note that using a pre-trained ViT considerably lowers the computational expense involved in training from scratch, while taking advantage of the rich feature representations obtained from large data sets. For the sake of reproducibility, we employed the ViT model available at the following URL: https://huggingface.co/docs/transformers/model_doc/vit#transformers.ViTForImageClassification.
Before feeding the ViT model with images, they are divided into a sequence of fixed-size non-overlapping patches. It is important to note that the transformer sequence length is inversely proportional to the square of the patch size. Consequently, while smaller patches allow for more detailed image analysis, they also increase computational complexity [41]. As a compromise between computational efficiency and array area covered, a patch size of $16\times 16$ pixels was selected, corresponding to an actual area of $25\,\rm m\times 25\,\rm m$.
@@
The model was trained on our problem data for 30 additional epochs using a batch size of 32, the default learning rate of $5\times 10^{-5}$, a warm-up ratio of $10\%$, and the mean square loss function. This pipeline was done with the Hugging Face Transformers open-source library, which provides APIs and tools to download and train state-of-the-art pre-trained models [55].
The input of the model is an image created by mapping the positions of the WCDs and assigning colors based on the total signal in each tank. Based on the optimal efficiency attained during the experimental stage, the Matplotlib [56] Magma colormap was utilised, with a dark background selected to depict the absence of energy in those areas. To highlight signal patterns within the footprint and cluster stations with comparable signals, a contour plot fills the space between the WCDs, using a grid with cells similar in size to the stations. This contour plot represents the logarithmic signal intensity levels, spaced over the range of observed signals, employing five equally spaced levels to illustrate gradients in signal distribution. The chosen number of levels balances the need to group stations effectively without over-segmenting the signals. Although fewer levels fail to capture meaningful groupings, higher numbers introduce unnecessary segmentation. The test with twenty levels yielded comparable results, supporting the suitability of five levels for this analysis. The continuous signal range of the WCDs and the discrete contour use an identical colormap to highlight stations with marginally higher or lower signals within a group. A logarithmic scale is applied to prevent saturation of the color scale due to the high signals in the stations near the shower core.
To reduce the noise of background or late-arriving particles, the shower plane is fitted using the positions of all active stations, with the $T_{0}$ of each station as the Z coordinate [57]. The $T_{0}$ for each station is defined as the timing bin within the signal trace where it records a minimum of 3 photoelectrons across all PMTs, with time bins of $1\,\rm ns$ [7]. It was found that with an average background of 23 stations, the plane fit successfully converges when at least 30 stations are present in the event. Therefore, only events with more than 30 active stations (signal + background) were considered for analysis. Following this procedure, only stations with a temporal offset of less than $50\,\rm ns$ from the fitted shower plane were selected. By applying the shower plane fit, approximately $40\%$ of the noise stations are excluded, while only about $20\%$ of the signal stations are eliminated. It should be noted that most of the excluded signal stations are those arriving late and not correlated with the shower plane, potentially lacking additional contribution to the footprint information. Thus, while identical results are observed for noise-free events, the approach effectively reduces the degradation caused by noise.
In addition, as part of the preprocessing steps, the images are rescaled to the same resolution ($224\times 224$) and normalised across the RGB channels with mean (0.5, 0.5, 0.5) and standard deviation (0.5, 0.5, 0.5) [51].
@@
Finally, the output of the ViT is defined as a classification probability, denoted as $P\in\left[0;1\right]$, that a proton has induced the analyzed shower.
@@
IV Results
This section assesses the gamma/hadron separation capability of the method by calculating the background rejection factor, $1-\varepsilon_{p}$, at a signal selection efficiency of $\varepsilon_{\gamma}\sim 80\%$. This value was chosen to ensure the identification of the majority of gamma-induced events. Moreover, the Appendix V.1 provides a description of the background rejection factor corresponding to the optimised gamma efficiency. This optimisation aims to maximise the Q-factor, defined as the ratio of signal selection to the square root of background rejection. A ViT model was trained for each energy bin and tested under various noise conditions, as detailed later.
Subsection E presents a realistic comparison of various gamma/hadron separators developed by the HAWC experiment, evaluated under the same experimental conditions.
For an EAS array with the given characteristics, the number of stations triggered by atmospheric muons within a 200200200\,200ns time window is expected to follow a Poisson distribution with a mean of 23 stations [57]. To assess the resilience of the method against noise from atmospheric muons, each shower event is contaminated with a random number of stations based on this distribution. The muonic signal added to the contaminated tanks is randomly selected using a probability function derived from the signal distribution of approximately 150 000150000150\,000150 000 single muon events from proton showers in the $\sim 1\,\rm TeV$ energy bin.
@@ IV.1 Resilience to Noise from Atmospheric Muons and Low-Energy Proton Showers
For each chosen reconstructed energy bin, we train and test a ViT model with both the original shower-derived images and additional images featuring contamination. This method considers the substantial variation in the shower footprint across different energy bins (see Figures 3a and 4a), which could otherwise lead to ambiguity without providing explicit energy information to the transformer. For instance, the footprint of low-energy gamma rays, often exhibiting dispersed signals and unclear cores, could be misclassified as hadronic showers. A unified model covering all energy bins would require integrating reconstructed energy into both training and inference, necessitating architectural modifications beyond the scope of this work. Instead, our focus remains on demonstrating the viability of the footprint-based technique across multiple energy bins, with an emphasis on the low-energy regime relevant to gamma-ray observatories.
Figure 5 presents the background rejection efficiency ($1-\varepsilon_{p}$) for the test data set, indicating that the addition of noise from low-energy proton showers has a minimal effect on the gamma/hadron separation technique, with only slight performance variations across the energy bins. However, as the energy increases, the degradation resulting from the overlap of both atmospheric muons and low-energy proton showers becomes more pronounced. This trend is expected to be related to the fact that low-energy showers rarely contain muons or signal clusters, which are similar to the noise, causing the model to rely less on those features for classification.
Overall, the method shows exceptional background rejection across all tested scenarios. This result demonstrates that there is enough information in the shower footprint to perform an outstanding gamma/hadron separation.
Figure 5: The gamma (top) and proton (bottom) selection efficiency at different energies and noise sources.
IV.2 Performance for different zenith angles
A dataset consisting of gamma-ray showers at a zenith angle of $\theta_{0}=30^{\circ}$ and protons within the range $\theta_{0}\in[25^{\circ};35^{\circ}]$ was simulated to assess the method’s effectiveness for inclined events. For a fair comparison, the simulations were conducted using the same methodology as with vertical events: the distribution of the shower core was homogeneous in the detector area and the energy reconstruction strategy used in Figures 1 and 2 was performed for this particular zenith angle bin.
The background noise is independent of the primary particle; thus, the same background noise and degradation are expected for other zenith angles and only events without contamination have been used for comparison. Since a large number of events are required to train these algorithms, the model trained with vertical events in Subsection IV.1 was employed for the comparison.
Figure 6 presents the proton rejection efficiency ($1-\varepsilon_{p}$) attained by ViT models for shower events with zenith angles of approximately $\theta_{0}\sim 10^{\circ}$ and $\theta_{0}\sim 30^{\circ}$. A similar level of degradation was observed across all energy ranges, indicating overall robust performance for inclined events. It should be mentioned that further improvements might be achievable by training the network with specific angle bins, analogous to the approach taken for energy. However, implementing an angular reconstruction method is beyond the scope of this study.
Figure 6: Dependence of the gamma (top) and proton (bottom) selection efficiency on the primary zenith angle at different energy bins for shower events without noise.
IV.3 Dependence on the position of the shower core
To evaluate how the performance of the gamma/hadron discrimination model is influenced by the position of the shower core, we examined the model’s output probabilities across different core positions, including those from off-array events.
To achieve this, the test shower events were iteratively simulated with their cores positioned at specific distances from the array center: $r_{\rm core}=\left\{0;50;100;150;170;180;190;200\right\}\,\rm m$. The models detailed in Subsection IV.1, trained exclusively on on-array events, were used for the evaluation.
The mean output probability of the model, shown in Figure 7, demonstrates that proton-induced showers consistently produce high probabilities close to 1, regardless of the position of the core. For gamma-induced showers, the model outputs probabilities close to 0 for on-array events, maintaining good separation from protons. However, for off-array events, the gamma probabilities increase, reducing the separation capability. Despite this, the results suggest that effective gamma/hadron discrimination remains achievable for events with core distances $r<180\,\rm{m}$, which is beyond the array radius of $160\,\rm{m}$. It is important to note that the footprint of the shower core at these energies may cover an area of a few tens of meters (see Figures 3 and 4), so most or all of it would be lost for off-array events. Although Figure 7 exclusively presents events in the $500\,\rm GeV$ energy bin, similar behaviour was also observed in other energy bins.
Figure 7: Average output for proton and gamma-induced showers with a reconstructed energy of roughly $500\,\rm GeV$ in bins of the shower core distance to the center of the detector array. The dashed line indicate the limit of the array.
IV.4 Dependence on the array fill factor
(a) $\rm FF=85\%$ (5 72057205\,7205 720 WCDs, original).
(b) $\rm FF=65\%$ (4 30643064\,3064 306 WCDs).
(c) $\rm FF=40\%$ (2 84628462\,8462 846 WCDs).
Figure 8: Footprint of a proton event with a reconstructed energy close to $1\,\rm TeV$ for various array configurations exhibiting different fill factors.
To ensure that the observed performance improvements were due to the proposed method rather than external factors such as the array layout or its fill factor, two additional configurations were developed. To minimise the need for repetitive and computationally expensive Geant4 simulations, these layouts were created by uniformly reducing the number of stations in the original array (see Figure 8). The original layout consisted of 5 72057205\,7205 720 WCDs arranged in a hexagonal pattern. The first alternative layout, with 4 30643064\,3064 306 WCDs and a fill factor of $65\%$, was created by removing the central station in each hexagonal grouping. The second layout featured a square arrangement of 2 84628462\,8462 846 WCDs, resulting in a fill factor of $40\%$. As this study focused on the sub-TeV energy range of a gamma-ray observatory, configurations with fill factors below $40\%$ were not explored.
Figure 9 shows the rejection efficiency achieved when using this method with the different array layouts. The models discussed in Subsection IV.1 were further trained for 10 additional epochs with a subset of training showers to accommodate the new fill factors. Both configurations suffer from a performance degradation, which is more noticible at the lower energy bin. However, the model remains robust and still provides excellent gamma/hadron separation with proton rejection factors comfortably below $2\cdot 10^{-2}$ with a high signal efficiency and $E_{r}\geq 500\,\rm GeV$.
Figure 9: The gamma (top) and proton (bottom) selection efficiency obtained using arrays with $\rm{Fill\,Factor}=\left\{40;65;85\right\}\%$.
IV.5 Comparison with other gamma/hadron discriminators
Although this method showed excellent performance compared to other observatories, this cannot be discerned from other factors such as the detector design, altitude at which the observatory is placed, or simulation setup, which can also influence the result and bias the comparison. For a fairer assessment, the gamma/hadron separators typically used by the HAWC observatory have been adapted for this specific detector design and computed for the same simulated events. Note that these discriminators originally used the PMT information due to the large size of the HAWC tanks; here, having smaller tanks, the signals at the station level were used to compute these variables. Moreover, this study did not incorporate geometry reconstruction; therefore, the Monte Carlo (MC) core position and angular measurements were employed to determine these variables. More details on the implementation of these discriminators can be found in the Appendix V.2. The following gamma/hadron separators have been implemented:
  • LDF $\chi^{2}$: a $\chi^{2}$ fit of the Lateral Distribution Function (LDF) using a modified Nishimura-Kamata-Greisen (NKG) function [58].
  • Parameter for Identifying Nuclear Cosmic rays (PINCness): a $\chi^{2}$-like measure of the smoothness of the charge footprint of the showers [59].
  • LIC: logarithm transformation of the inverse of the compactness parameter, an empirical parameter originally developed by the Milagro Collaboration [59, 60].
Figure 10: The gamma (top) and proton (bottom) selection efficiency obtained by the HAWC gamma/hadron discriminators in showers contaminated by noise atmospheric muons and low-energy proton showers. The results integrate the shower plane fit to mitigate noise effects and are contrasted with the results reported by HAWC at a lower altitude (represented by the gray dotted line) [58].
Figure 10 presents the proton rejection efficiency ($1-\varepsilon_{p}$) achieved by the gamma/hadron separators, alongside a comparison with the footprint-based method and the most recent results published by the HAWC Collaboration [58], which combines LDF $\chi^{2}$ and compactness. Showers with noise from both atmospheric muons and low-energy proton showers have been used to ensure a realistic comparison (the same comparison using events without noise can be found in Figure 13). A combination of the other three discriminators that use a Boost Decision Tree (BDT) yielded significant enhancements compared to using single variables. The BDT was trained using a mean squared error loss and configured with $100$ estimators, a maximum depth of 6, and a learning rate of $0.3$. Our analysis of the alternative discriminators demonstrated improved performance relative to the results reported by HAWC. This discrepancy could be attributed to the optimistic computation of these variables using the MC core and angle values, the variation in altitude –our simulated observatory is located at 5 $5\,200\,\rm m$ a.s.l, whereas HAWC is situated at approximately 4 $4\,100\,\rm m$ a.s.l– or the differences in the detector design.
These results also demonstrate that, under the same experimental conditions, the footprint-based method performs better than the other gamma/hadron separators, especially at the GeV energy range. It is important to note that although the plane fit helped reduce background noise, it did not completely remove it. The footprint-based method showed better resilience to noise than the other discriminators at equal conditions. For instance, the LDF $\chi^{2}$ parameter showed an excellent performance for events without noise (see Figure 13), being the best among the HAWC g/h discriminators, but noise can affect the LDF fits, causing greater degradation in classical approaches.
IV.6 Dependence on the muon content of the shower